Robert recorded the number of calls he made at work during the week:

Day Calls
Monday 20
Tuesday 12
Wednesday 10
Thursday 18


He expected to make 15 calls each day. To determine whether the number of calls follows a uniform distribution, a chi-square test for goodness of fit should be performed (alpha = 0.05).



Using the data above, what is the chi-square test statistic? Answer choices are rounded to the hundredths place.

  • a.)

    0.67

  • b.)

    0.42

  • c.)

    4.54

  • d.)

    3.75

SUBMIT MY ANSWER
Answer #1

Answer

The value of Chi-Squared statistic = 4.54

hence, option c) is correct.

Solution:

The following table is obtained:

Days Calls (Observed) Expected (fo-fe)2/fe
Mon 20 60*0.25=15 (20-15)2/15 = 1.67
Tue 12 60*0.25=15 (12-15)2/15 = 0.6
Wed 10 60*0.25=15 (10-15)2/15 = 1.67
Thurs 18 60*0.25=15 (18-15)2/15 = 0.6
\sum= 60 60 4.54

where,

fo = observed frrequency = Oi

fe = expected frequency = Ei

Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

H_0: p_1 = 0.25, p_2 = 0.25, p_3 = 0.25, p_4 = 0.25

Ha​: Some of the population proportions differ from the values stated in the null hypothesis

This corresponds to a Chi-Square test for Goodness of Fit.

Rejection Region

Based on the information provided, the significance level is

α=0.05,

the number of degrees of freedom is df = 4 - 1 = 3

so the rejection region for this test is

R = { χ2 : χ2 > 7.815}

Test Statistics

The Chi-Squared statistic is,

\chi^2 = \sum_{i=1}^n \frac{(O_i-E_i)^2}{E_i} = 1.67+0.6+1.67+0.6 = 4.54

Decision about the null hypothesis

Since it is observed that

\chi^2 = 4.533 \leq \chi_c^2 = 7.815

it is then concluded that the null hypothesis is not rejected.

Conclusion

It is concluded that the null hypothesis Ho is not rejected.

Therefore, there is NOT enough evidence to claim that some of the population proportions differ from those stated in the null hypothesis, at α=0.05 significance level.